### 14 Mar 2026 v1

# Is Seeing Believing? Evaluating Human Sensitivity to Synthetic Video

Abstract
Advances in machine learning have enabled the creation of realistic synthetic videos known as deepfakes. As deepfakes proliferate, concerns about rapid spread of disinformation and manipulation of public perception are mounting. Despite the alarming implications, our understanding of how individuals perceive synthetic media remains limited, obstructing the development of effective mitigation strategies. This paper aims to narrow this gap by investigating human responses to visual and auditory distortions of videos and deepfake-generated visuals and narration. In two between-subjects experiments, we study whether audio–visual distortions affect cognitive processing, such as subjective credibility assessment and objective learning outcomes. A third study reveals that artifacts from deepfakes influence credibility. The three studies show that video distortions and deepfake artifacts can reduce credibility. Our research contributes to the ongoing exploration of the cognitive processes involved in the evaluation and perception of synthetic videos, and underscores the need for further theory development concerning deepfake exposure.
Keywords Deepfakes $\cdot$ misinformation $\cdot$ social media $\cdot$ synthetic media $\cdot$ disfluency $\cdot$ cognitive psychology
1 Introduction
Recent advances in machine learning architectures capable of synthesizing highly realistic video and audio material, so-called deepfakes, threaten the integrity of videos as reliable information sources Westerlund (2019). A central concern is that malicious actors can use the technology to create and spread false evidence, as deepfakes enable the creation of convincing impersonations of influential and trusted figures Köbis et al. (2021); Mink et al. (2024), and are frequently deployed in contexts of sexual harassment Wolfe et al. (2023); Gebru and Denton (2024). The introduction of deepfake videos raises concerns as people seem to be particularly susceptible to video-based disinformation Sundar et al. (2021); Lee and Shin (2022). Despite the growing awareness of issues concerning deepfakes, we lack a general understanding of how people perceive and evaluate synthetically generated media Tahir et al. (2021); Gamage et al. (2022).
The visual aspect of a video may influence how its message is perceived in several ways: It may contain arguments, focus or distract the audience’s attention, and serve as cue for heuristic evaluations Peng et al. (2023). A key heuristic in processing visual information is processing fluency, the subjective ease of processing a stimulus. When information is easier (or harder) to process than expected, this metacognitive “feel” systematically shifts judgments such as truth, trust, and liking Whittlesea and Williams (2000). Prior research has found evidence for the influence of processing fluency on judgments of credibility and trust in numerous media types Schwarz et al. (2021). The legibility of text Reber and Schwarz (1999), the quality of audio recordings Newman and Schwarz (2018), and the presence of nonprobative photos alongside a statement Newman et al. (2012) were found to affect how credible people deemed information to be. Nevertheless, the effects of processing fluency of videos remain mostly underexplored.
Given the unique characteristics of videos and the increasing prevalence of deepfakes, studying the effects of visual and auditory video aspects on credibility becomes crucial Gamage et al. (2022). This is because artifacts remain the major indicator of artificially generated content, despite the increasing quality of deepfake videos Ciftci et al. (2020). These artifacts often manifest as unnatural movements, extraneous limbs, inconsistencies in the subject’s lighting, and similar features that do not occur in recorded videos Matern et al. (2019).
Distortions and deepfake artifacts can reduce processing fluency because they violate viewers’ expectations about how recorded videos typically look and sound. Such low-level irregularities act as perceptual expectancy violations that make the video subjectively harder to process, and may therefore serve as heuristic cues in credibility judgments. This is especially relevant in online video environments, where high information density and limited attention promote intuitive, heuristic modes of evaluation Greifeneder et al. (2011); Lang (2000). Under these conditions, which are common on platforms such as YouTube and TikTok Metzger et al. (2010); Sundar et al. (2021), viewers are likely to rely on directly accessible perceptual cues rather than analytical assessment. We therefore examine whether visual distortions, auditory distortions, and deepfake artifacts in faces and voices systematically affect video credibility, processing fluency, and related evaluation outcomes.
For HCI, credibility and authenticity are interactional judgments shaped by interface cues as much as by content. Online videos are typically encountered in environments that impose cognitive load and encourage heuristic processing Lang (2000); Metzger and Flanagin (2015), meaning that subtle audiovisual irregularities become salient inputs to users’ credibility assessments. Yet, we lack empirical evidence on how users in realistic viewing conditions respond to artifacts in synthetic media Hilligoss and Rieh (2008); Sundar (2008); Metzger and Flanagin (2015).
In this paper, we address three questions: (1) do visual distortions reduce video credibility; (2) do auditory distortions and audio–visual timing irregularities have similar effects; and (3) do deepfake artifacts in faces and voices, separately and combined, undermine credibility and learning? Across three preregistered online experiments we find that visual distortions and deepfake artifacts reliably lower message credibility and source vividness, while leaving learning largely unaffected.
We find that deepfake artifacts reduce both message credibility and perceived source vividness, while processing fluency remains unaffected. Our contributions are: (1) empirical evidence that deepfake artifacts undermine credibility in subtle but measurable ways; (2) an analysis of how distortions and deepfake artifacts in audio and video shape user judgments; and (3) open and reproducible research resources, including all data, analyses, and stimuli shared on OSF https://doi.org/10.17605/OSF.IO/T3AWB). Together, these findings advance our understanding of human sensitivity to deepfakes and provide a basis for developing theory, policy, and interventions for deepfake exposure.o
2 Background and related work
2.1 Deepfakes and Artifacts
Deepfakes are videos in which faces and/or voices have been synthetically generated using deep-learning models. The technology enables convincing impersonations with comparatively little expertise Kaate et al. (2023); Roose (2018) and is increasingly used in deceptive contexts Dobber et al. (2021). For the present work we distinguish three categories: (1) recorded videos, captured without AI synthesis; (2) distorted videos, where visual or auditory degradations are added without using generative models; and (3) deepfake videos, in which the facial appearance and/or narration have been synthesized.
Deepfake generation pipelines introduce characteristic artifacts because models must estimate facial geometry, motion, illumination, and texture, and temporally align synthesized audio and video. Imperfections in these steps produce residual cues that differentiate synthetic from recorded footage, such as subtle inconsistencies in shading, boundary alignment, facial dynamics, or mouth–speech synchrony Matern et al. (2019); Li et al. (2020); Yang et al. (2019); Ciftci et al. (2020). While some artifacts may be imperceptible to viewers but detectable algorithmically Agarwal et al. (2020), others manifest as visible irregularities or atypical movement patterns.
Despite these cues, viewers often struggle to reliably distinguish deepfakes from recorded videos. Prior studies show that many people fail to identify deepfakes even when aware of their existence Korshunov and Marcel (2021); Groh et al. (2022); Köbis et al. (2021); Mink et al. (2024), and deepfake political videos frequently receive moderate credibility ratings Lee and Shin (2022); Hwang et al. (2021). Uncertainty is widespread, and judgments are easily confounded by factors such as familiarity, liking, or sympathy for the depicted individual Vaccari and Chadwick (2020); Köbis et al. (2021). This difficulty underscores the need to understand how specific kinds of visual and auditory artifacts shape users’ evaluations.
2.2 Heuristics and processing fluency in credibility judgments
Credibility refers to an individual’s judgment of the veracity of the content of communication Appelman and Sundar (2016). Online environments complicate such judgments: information volume is high, source cues are sparse or ambiguous, and professional verification is limited Callister (2000); Metzger (2007); Metzger and Flanagin (2015); Sohrawardi et al. (2024). Under these conditions of overload and limited cognitive resources, people rely on heuristics; that is, readily accessible cues that can be evaluated quickly but are fallible Lang (2000); Hilligoss and Rieh (2008). Social indicators such as endorsements, comments, or source reputation often serve this function Metzger et al. (2010), and popularity has been shown to be influential for credibility assessments explicitly in the context of deepfake content Jin et al. (2025).
Deepfakes complicate credibility judgments further because they can appropriate trusted identities and obscure more diagnostic cues. Although synthetic videos may appear plausible, they often contain subtle irregularities. Such unexpected visual or auditory features create expectancy violations; deviations from what viewers assume recorded videos should look or sound like, which can reduce processing fluency, the subjective ease of processing a stimulus Whittlesea and Williams (2000). Processing fluency reliably influences credibility judgments across media: less legible text Reber and Schwarz (1999) and lower-quality audio Newman and Schwarz (2018) reduce perceived truthfulness Schwarz et al. (2021).
People particularly rely on fluency when external knowledge is limited Alter and Oppenheimer (2009). One proposed mechanism is familiarity: fluently processed information feels more familiar and is therefore judged as more likely to be true Dechêne et al. (2010); Whittlesea et al. (1990). In the context of deepfakes, this suggests that viewers may turn to the “feel” of ease or difficulty when evaluating unfamiliar videos or presenters. However, it remains unclear whether and how such fluency-based heuristics operate for videos, and especially for deepfake-generated faces and voices, where multiple perceptual channels may produce competing or ambiguous cues.
2.3 Credibility assessment as an interaction problem
In online environments, credibility judgments are shaped by an interactional process that includes the viewer and the design of digital interfaces. Users evaluate authenticity within feeds, video players, and recommendation systems that structure attention and impose cognitive constraints Hilligoss and Rieh (2008); Metzger and Flanagin (2015). HCI research has shown that interface-level cues such as modality, visual clarity, and system agency systematically influence heuristic credibility judgments Sundar (2008); Sun et al. (2024); Zieglmeier and Lehene (2021). In this perspective, deepfake artifacts can be understood as perceptual cues embedded in the interaction environment: they are surfaced to the user through the interface in the same way as other credibility-relevant signals. Studying how users respond to these artifacts therefore addresses the central question of how interactive systems mediate trust and uncertainty when synthetic media are encountered in real-world interface contexts.
3 Studies I–III
Across three preregistered online experiments, we systematically examined how manipulations to visual and auditory properties of an educational video, including deepfake artifacts, affect processing fluency, credibility, and learning, in addition to select metacognitive outcomes. Study I tested the hypothesis that reducing the processing fluency of video imagery makes it less credible. Study II extended this approach to speech clarity and audio–visual synchrony, and added learning and additional metacognitive outcomes. Studies I and II showed that visual distortions affect perceived processing fluency and credibility. Study III extended this progression to ecologically realistic deepfake artifacts by independently manipulating synthesized faces and synthesized narration.
4 Methods
4.1 Design
All three studies employed between-subjects experimental designs with a single independent variable. Participants were randomly assigned to one condition that differed in the presentation style of an educational video.
Study I used a two-level design, comparing a baseline condition with a reduced visual clarity condition. Study I tested the hypothesis:
  1. H1
    Reducing the processing fluency of a video with visual distortions makes it less credible.
Study II used a four-level design. Participants were randomly assigned to reduced visual clarity, reduced speech clarity, audio–visual asynchrony, or baseline. The study tested three hypotheses:
  1. H1
    Reducing the processing fluency of video with visual distortions makes it less credible (replicating Study I).
  2. H2
    Reducing the processing fluency of the speech in a video with auditory distortions makes it less credible.
  3. H3
    Asynchrony between video and audio reduces the video’s processing fluency and makes it less credible.
Study III used a four-level design. Participants were randomly assigned to synthesized video, synthesized narration, synthesized video and narration, or baseline. The study tested three hypotheses:
  1. H1
    A video with deepfake-generated visuals is processed less fluently than a recorded video and is less credible.
  2. H2
    A video with deepfake-generated narration is processed less fluently than one with recorded narration and is less credible.
  3. H3
    A video with deepfake-generated visuals and narration is processed less fluently than one with recorded narration and is less credible.
4.2 Preregistration
All three studies were preregistered on the Open Science Framework along with their hypotheses, study plans, and statistical analysis plans (see Supplementary Materials). All methods were carried out in accordance with the relevant guidelines and regulations for empirical research, and the experimental protocols were approved by ANONYMIZED.
4.3 Participants
All participants were US-based, English-speaking adults recruited via the online platformProlific [^1] https://prolific.com.
Study I
Three hundred and sixty US Americans were recruited in July 2023. After removal of $27$ participants not adhering to the preregistered quality criteria, data from $333$ participants were analyzed. The cohort comprised $174$ males, $147$ females, and $12$ participants who identified with other genders. Participants’ mean age was $40.6$ years ($SD=13.2$). Educational levels ranged from ‘High school or less’ ($n=139$), ‘Bachelor’ ($n=130$), ‘Master’ ($n=56$), to ‘PhD’ ($n=8$). The mean completion time was about four minutes ($M=$ $234\text{\,}\mathrm{s}$,SD=114,\ SD=114).
Study II
In May 2023, we recruited US-based English-speaking participants via Prolific. A pilot study with 16 participants indicated an effect size of Cohen’s $f=0.22$. We recruited $400$ participants (see Appendix 1 for power analysis details). After exclusion of participants not meeting the preregistered quality criteria, data from $355$ participants were analyzed. The cohort comprised $187$ males, $156$ females, and $12$ participants who identified with other genders. Mean age was $43.0$ years ($SD=12.9$). Educational levels ranged from ‘High school or less’ ($n=150$), ‘Bachelor’ ($n=154$), ‘Master’ ($n=41$), to ‘PhD’ ($n=10$). The mean completion time was slightly above 15 minutes ($M=$ $980\text{\,}\mathrm{s}$,SD=365,\ SD=365).
Study III
In September 2023, we recruited $400$ US-based English-speaking participants via Prolific (see Appendix 1 for power analysis). After exclusion of participants not meeting the required quality criteria, data from 365 participants were analyzed. The cohort comprised 185 males, 170 females, and 10 participants who did not self-describe as either of those. Mean age was $44.4$ years ($SD=16.5$). Educational levels ranged from ‘High school or less’ ($n=129$), ‘Bachelor’ ($n=177$), ‘Master’ ($n=47$), to ‘PhD’ ($n=12$). The mean completion time was about 17 minutes ($M=$ $1005\text{\,}\mathrm{s}$,SD=363,\ SD=363).
4.4 Dependent variables
Across the three studies, we used self-report measures and multiple-choice knowledge tests; Study I employed a subset of the measures used in Studies II and III (see details in Table S1 and  Table S2).
Processing fluency
Following Graf et al. (2018), subjective processing fluency was measured with a slider, with internally stored responses between 0 (difficult) and 100 (easy). This measure was used in all three studies.
Message credibility
Based on the scale by Appelman and Sundar (2016), message credibility was computed as the mean of three 7-point Likert items accurate, authentic, and believable, with response options ranging from 1 (“describes very poorly”) to 7 (“describes very well”). This scale was used in all three studies.
Digital alteration detection
In all three studies, digital alteration detection was measured using the question “Do you think the video has been digitally altered?” with response options “Yes”, “No”, and “I don’t know”.
Learning
In Study II and Study III, conceptual and factual learning were measured as the pre-to-post difference in a multiple choice questionnaire, following guidelines by Petersen et al. (2021). Pre- and post-stimulus knowledge questionnaires were administered with mixed item order and embedded attention checks.
Source vividness, attentional focus, and liking
In Study II and Study III, source vividness and attentional focus were measured on 7-point Likert scales anchored at “strongly disagree” and “strongly agree”. Message credibility (as above) and liking were measured as 7-point Likert scales anchored between “describes very poorly” and “describes very well”. For all multi-item scales, we report means.
4.5 Stimuli
Common stimulus video
An educational video was produced to serve as a common stimulus source for all three studies. The video was $447\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}\mathrm{o}\mathrm{n}\mathrm{d}\mathrm{s}$ long and featured an actor presenting a script adapted from Petersen et al. (2021). All video editing was performed in Adobe After Effects 23.2.1 (see Appendix 2 and Appendix 3 for details on video processing). The original script consisted of sections on general virology, measles, Zika, and COVID-19; we excluded the COVID-19 section to avoid interference from prevalent, non-scientific interpretations of the disease.
We chose the educational-video genre for four reasons. First, it allowed us to present complex, domain-specific information that is likely unfamiliar to most participants, thereby reducing the likelihood that credibility judgments could be resolved by simply validating content against prior knowledge. Second, the topic and presentation were designed to be ideologically neutral, limiting the influence of pre-existing attitudes or motivated reasoning on credibility assessments. Third, the format is familiar and ecologically valid for contemporary online platforms (e.g., YouTube, TikTok), where viewers routinely encounter informational videos. Finally, we used an unknown presenter against a neutral background to minimize extraneous cues such as source reputation, studio branding, or contextual set dressing that could independently affect credibility.
Study I: Visual fluency manipulation
In Study I, we used the video’s first section on general virology ($118$ seconds) as the source for the stimuli. In the baseline condition, the video was left unaltered. In the reduced visual clarity condition, the video featured a ghosting filter that superimposed a semitransparent duplicate of the original video over the original (see 1(b)). The speed of the superimposed video was manipulated to transition gradually between leading and lagging. These manipulations were chosen to distort the viewing experience without introducing elements that might be perceived as deceptive, thereby reducing the likelihood that credibility ratings were driven merely by suspicion.
Study II: Visual, speech, and synchrony manipulations
In Study II, the full $447\text{\,}\mathrm{s}\mathrm{e}\mathrm{c}\mathrm{o}\mathrm{n}\mathrm{d}\mathrm{s}$ video was used unchanged in the baseline condition and as the basis for all three treatment stimuli. As in Study I, the reduced visual clarity condition featured a ghosting filter. Here, the “ghost” layer was only $30\%$ opaque, to achieve a less distorted viewing experience.
The reduced speech clarity condition applied a mild echo to the original narration. We selected this transformation to lower listening fluency while preserving all lexical content and timing, thereby minimizing the risk that credibility judgments were driven by suspicions of deliberate tampering rather than by audio quality per se (see Appendix 3 for details). The audio–visual asynchrony stimulus featured a one-second delayed audio track that brought the speech and the presenter’s lip movements out of alignment.
Study III: Deepfake-based manipulations
In Study III, the same recorded educational video served as the baseline stimulus. Deepfake versions of the presenter’s face and narration were created to produce the three experimental conditions (see Figure 2).
Figure 1: Stills of the two stimulus videos used in Study I illustrating the differences between the experimental conditions: (a) a still from the unaltered, recorded video shown to participants in the baseline condition, and (b) a still from the distorted video shown to participants in the reduced visual clarity condition
Figure 2: Stills showing the stimulus videos used in Study III illustrating the visual differences between the experimental conditions: (a) a still from the unaltered, recorded video shown to participants in the conditions baseline and synthesized narration, and (b) a still from the deepfake video shown to participants in the synthesized video and synthesized video and narration conditions.
We used the Latent Image Animator with “vox-pt” weights by Wang and colleagues Wang et al. (2022) to synthesize a deepfake version of the video. This approach requires a $256\times 256$ close-up video of a face and a $256\times 256$ portrait picture to generate a $256\times 256$ video of the portrait picture exhibiting the behavior captured in the video. The resulting animated portrait was then pasted into and aligned with the source video to obtain the deepfaked stimulus video. We synthesized a version of the presenter’s narration by fine-tuning a So-Vits-Svc-fork model [^2] https://github.com/voicepaw/so-vits-svc-fork. The model was trained on a $14.25\text{\,}\mathrm{m}\mathrm{i}\mathrm{n}\mathrm{u}\mathrm{t}\mathrm{e}\mathrm{s}$ long recording of a diverse set of English sentences adapted from the Harvard sentences [^3] https://harvardsentences.com. Once trained and provided with the audio of the source video, the model synthesized a version of the original narration in the new voice. The synthesized video stimulus was created by replacing the presenter’s face in the source video with the deepfake animation while retaining the original audio. The synthesized narration stimulus replaced the source video’s original narration track with the synthetic version. For the synthesized video and narration condition, we combined the deepfake stimulus video with the synthesized narration track.
4.6 Procedure
All studies were administered online using Gorilla. In all three experiments, participants provided informed consent, confirmed audio readiness, reported demographics, and affirmed English proficiency. They then viewed their assigned stimulus video with playback controls disabled and completed attention checks. After the video, participants responded to the digital-alteration item and were debriefed.
Study I
Following video exposure, participants completed the processing-fluency and message-credibility measures (randomized order).
Studies II and III
Studies II and III followed the same core procedure but additionally included pre- and post-stimulus knowledge tests with mixed item order and attention checks. After viewing the video, participants completed processing-fluency and message-credibility measures as well as scales for source vividness, liking, and attentional focus (randomized order).
5 Results
Because most measures were non-normally distributed, we deviated from the preregistered plan in using conservative non-parametric tests (Kruskal–Wallis; Mann–Whitney $U$ for post hoc comparisons with Bonferroni correction where indicated). Effect sizes are reported as rank-biserial correlations. As preregistered for Study III, Bayesian estimation complements the frequentist tests (see Appendix for full details).
5.1 Study I
Message credibility
Message credibility ratings (1–7) were higher in the baseline condition (M=6.2,SD=0.8M=6.2,\ SD=0.8) than in the reduced visual clarity condition (M=5.8,SD=1.3M=5.8,\ SD=1.3; see Figure 3). The Mann–Whitney $U$ test indicated a significant difference between the conditions, $p<0.0001$. The rank-biserial correlation was $0.50$.
Processing fluency
Processing fluency ratings averaged $86.7$ ($SD=19.0$) in the baseline condition and $57.2$ ($SD=34.2$) in the reduced visual clarity condition (see Figure 3). The Mann–Whitney $U$ test indicated a significant difference between conditions, $p=0.004$. The rank-biserial correlation was $0.18$.
Figure 3: Measurements of Message Credibility (a) and Processing Fluency (b) in Study I divided by experimental condition. Dots represent mean values, horizontal lines indicate bootstrapped confidence intervals. Significant differences are indicated with brackets and asterisks: One asterisk (*) $p<0.05$, two asterisks (**) $p<0.01$, three asterisks (***) $p<0.0001$.
Alteration detection
More participants detected a digital alteration in the reduced visual clarity condition ($n=126$) than in the baseline condition ($n=22$). More participants were uncertain about digital alterations in the baseline group ($n=59$) than in the reduced visual clarity group ($n=32$). Similarly, more participants in the baseline group did not detect an alteration ($n=81$) compared to the reduced visual clarity group ($n=13$; see Figure 4). Building on work by Altay and Gilardi (2024), who found that the belief of text being AI-generated negatively impacting its perceived credibility, we expanded on our preregistered analysis and examined whether believing that the video was digitally altered was associated with lower credibility. A Spearman rank correlation indicated a moderate negative association between alteration detection and credibility (r(331)=.25,p<.001r(331)=-.25,\,p<.001), see Appendix 7 for details.
Figure 4: Percentage of responses to the alteration detection question by treatment group in Study I.
Exploratory analyses
Exploratory correlations indicated no significant relationship between age and processing fluency ($r(331)=0.06$, $p=0.26$) but a small positive correlation between age and message credibility ($r(331)=0.18$, $p=0.0008$). Kruskal–Wallis tests did not show significant relationship between participant education and processing fluency ($p=0.10$) or message credibility ($p=0.76$).
5.2 Study II
Message credibility
A Kruskal–Wallis test demonstrated a significant difference in message credibility scores between conditions, $H(3)=13.39$, $p=0.004$. Subsequent pairwise Mann–Whitney $U$ tests with Bonferroni adjustment revealed significant differences between the reduced speech clarity condition (M=6.22,SD=1.05M=6.22,\ SD=1.05) and the reduced visual clarity condition (M=5.87,SD=1.01M=5.87,\ SD=1.01); $p=0.024$, rank-biserial correlation =0.24=0.24, as well as between reduced speech clarity and audio–visual asynchrony (M=5.81,SD=1.23M=5.81,\ SD=1.23); $p=0.019$, rank-biserial correlation =0.25=0.25 (see Figure 5).
Processing fluency
A Kruskal–Wallis test demonstrated a significant overall difference in processing fluency scores between conditions, $H(3)=9.36$, $p=0.02$. However, the pairwise differences between conditions were not significant after Bonferroni correction (see Figure 5).
Learning
Kruskal–Wallis tests did not reveal significant differences in factual learning ($H(3)=5.65$, $p=0.13$) or conceptual learning ($H(3)=2.03$, $p=0.57$) between conditions (see Figure 5).
Figure 5: Measurements of Message Credibility (a), Processing Fluency (b), Factual and conceptual learning (c), and Source Vividness (d) in Study II divided by experimental condition. Dots represent mean values, horizontal lines indicate bootstrapped confidence intervals. Significant differences according to Bonferroni-corrected tests are indicated with brackets and asterisks: One asterisk (*) $p<0.05$, two asterisks (**) $p<0.01$.
Source vividness
A Kruskal–Wallis test indicated a statistically significant difference in source vividness scores across conditions, H(3)=17.12,p=0.001H(3)=17.12,\ p=0.001. Post hoc Mann–Whitney $U$ tests with Bonferroni adjustment showed significant differences between baseline (M=5.32,SD=1.31M=5.32,\ SD=1.31) and reduced visual clarity (M=4.65,SD=1.44M=4.65,\ SD=1.44); $p=0.01$, rank-biserial correlation =0.26=0.26, and between baseline and audio–visual asynchrony (M=4.62,SD=1.45M=4.62,\ SD=1.45); $p=0.01$, rank-biserial correlation =0.28=0.28. A significant difference was also observed between audio–visual asynchrony and reduced speech clarity (M=5.21,SD=1.32M=5.21,\ SD=1.32); $p=0.03$, rank-biserial correlation =0.24=0.24 (see Figure 5).
Liking
A Kruskal–Wallis test indicated a significant difference in liking across conditions, $H(3)=10.22$, $p=0.02$. Post hoc Mann–Whitney $U$ tests with Bonferroni adjustment revealed a significant difference between baseline (M=4.23,SD=0.94M=4.23,\ SD=0.94) and audio–visual asynchrony (M=3.83,SD=0.89M=3.83,\ SD=0.89); $p=0.03$, rank-biserial correlation =0.24=0.24.
Attentional focus
A Kruskal–Wallis test did not uncover a significant difference in attentional focus across conditions, $H(3)=6.37$, $p=0.09$.
Alteration detection
All groups included participants who selected “yes” when asked whether the video was digitally altered. The share of “yes” responses was largest in the reduced visual clarity condition ($63\%$), followed by audio–visual asynchrony ($41\%$), reduced speech clarity ($23\%$), and baseline ($10\%$). The proportion answering “I don’t know” was largest in the baseline condition ($43\%$), followed by audio–visual asynchrony ($37\%$), reduced speech clarity ($33\%$), and reduced visual clarity ($23\%$). This pattern and the lack of a substantial difference between baseline and manipulated conditions suggest generally high uncertainty regarding perceived video alteration that is most effectively resolved by overt visual alteration, but that persists despite the absence of any alteration (see Figure 6). Here too, a Spearman correlation indicated a negative association between alteration detection and credibility rating (r(353)=.27,p<.001r(353)=-.27,\,p<.001).
Figure 6: Percentage of responses to the alteration detection question in Study II by treatment group.
5.3 Study III
Message credibility
The manipulations significantly affected message credibility, $H(3)=14.95$, $p=0.002$. Post hoc Mann–Whitney $U$ tests with Bonferroni correction indicated a significant difference between baseline (M=6.18,SD=0.99M=6.18,\ SD=0.99) and synthesized video and narration (M=5.64,SD=1.17M=5.64,\ SD=1.17); $p=0.02$, rank-biserial correlation =0.24=0.24 (see Figure 8). Bayesian estimation corroborated this effect: the distribution of credible differences in message credibility between baseline and synthesized video and narration was entirely above 0 and almost entirely outside a region of practical equivalence corresponding to Cohen’s $d<0.1$ (see Figure 7).
Figure 7: Distribution of credible differences between the baseline video and the fully synthesized video. (a) shows the distribution of message credibility differences, (b) the distribution for processing fluency in Study III. Bayesian estimation Kruschke (2013) was used to calculate distributions of mean processing fluency and message credibility values for the baseline and synthesized video and narration conditions. The differences between the means are expressed in fractions of the pooled standard deviation (i.e., in units comparable to Cohen’s $d$). The 95% HDIs (High Density Intervals) indicate where the bulk of the credible differences fall. For processing fluency (b), the HDI is close to or includes 0, consistent with no meaningful difference between conditions. For message credibility (a), the HDI lies substantially above 0, and essentially all credible effect sizes exceed a small-effect threshold (Cohen’s $d=0.10$), indicating that the distance between the means is too large to accept $H_{0}$.
Processing fluency
The Kruskal–Wallis test indicated no statistically significant difference in processing fluency scores across conditions, $H(3)=2.35$, $p=0.503$ (see Figure 8). The Bayesian analysis yielded a 95% highest density interval (HDI) for the difference between baseline and synthesized video and narration that included 0 and lay largely within the region of practical equivalence, supporting the conclusion that $H_{0}$ can be accepted for processing fluency (see Figure 7).
Learning
Kruskal–Wallis tests indicated no significant differences in factual learning ($H(3)=0.72$, $p=0.87$) or conceptual learning ($H(3)=2.55$, $p=0.47$) between conditions (see Figure 8). Bayesian estimation similarly indicated no practically meaningful differences in learning scores between the baseline and experimental conditions on either subscale.
Figure 8: Measurements of Message Credibility (a), Processing Fluency (b), Learning (factual and conceptual combined) (c), and Source Vividness (d) in Study III divided by experimental condition. Dots represent mean values, horizontal lines indicate bootstrapped confidence intervals. Significant differences according to Bonferroni-corrected tests are indicated with brackets and asterisks: One asterisk (*): $p<0.05$, two asterisks (**): $p<0.01$.
Source vividness
The manipulations significantly affected source vividness, $H(3)=18.70$, $p=0.0003$. Post hoc Mann–Whitney $U$ tests with Bonferroni correction showed significant differences between baseline (M=5.23,SD=1.29M=5.23,\ SD=1.29) and synthesized video and narration ($p=0.007$, rank-biserial correlation =0.28=0.28), and between baseline and synthesized video ($p=0.002$, rank-biserial correlation =0.31=0.31; see Figure 8).
Liking
A Kruskal–Wallis test indicated no statistically significant differences in liking scores across conditions, $H(3)=5.30$, $p=0.15$.
Attentional focus
A Kruskal–Wallis test indicated no statistically significant differences in attentional focus scores across conditions, $H(3)=1.64$, $p=0.65$.
Alteration detection
All groups included participants who selected “yes” when asked whether the video was digitally altered. The share of “yes” responses was highest in the synthesized video and narration condition ($54\%$), followed by synthesized video ($42\%$), synthesized narration ($32\%$), and baseline ($14\%$). The proportion answering “I don’t know” was similar in the baseline ($39\%$) and synthesized narration ($38\%$) conditions, and lower in synthesized video ($31\%$) and synthesized video and narration ($28\%$). The synthesized video and narration condition featured the fewest participants selecting “no” ($18\%$), and the baseline condition the highest ($47\%$). Thus, even in the absence of synthetic content, more than $10\%$ of participants reported detecting an alteration and roughly a third were uncertain. This uncertainty is lower in conditions with synthetic video, which also feature the highest rate of indicated detections. The effects of synthetic narration alone are comparatively limited (see Figure 9). Among the three studies the negative association between alteration detection and credibility rating was strongest here (r(363)=.40,p<.001r(363)=-.40,\,p<.001)
Figure 9: Percentage of responses to the alteration detection question in Study III by treatment group.
6 Discussion
Our study examined how visual and auditory manipulations affect perceived video credibility in the context of deepfakes. As summarized in Table 1, the three studies form a cumulative progression from simple distortions to realistic deepfake artifacts. The results indicate that visual manipulations have a small but significant impact, though this effect cannot be explained by processing fluency.
Video Credibility
Our investigation into visual and auditory distortions and deepfake artifacts on video credibility across three studies reveals their significant influence on viewer perceptions.
In Study II, visual distortions and audio-video asynchrony reduced credibility compared to audio distortions alone but not the baseline, highlighting the role of sensory cues in credibility assessment. Study III found that only the combination of deepfake visuals and narration significantly reduced credibility compared to the baseline, suggesting a small additive effect. Contrary to Newman and Schwarz Newman and Schwarz (2018), we found no significant impact of audio manipulations on credibility.
Processing Fluency
Our findings do not conclusively demonstrate the role of processing fluency in video perception. Study I showed a correlation between processing fluency and credibility; this was not evident in Study II and III. The data showed participants responded in a dichotomous way; participants positioned the indicator either towards 0 (left) or 100 (right) from its default position in the middle of the range.
Alteration Detection
Groups of participants that indicated having and not having detected an alteration were present across all three studies, regardless of the presence of an manipulation. The data reveals a significant rate of inaccuracy in identifying distortions and deepfake videos, reflected in the high proportion of participants selecting “I don’t know.” These findings are consistent with Vaccari and Chadwick’s research, which links deepfakes to increased uncertainty Vaccari and Chadwick (2020). Interestingly, the baseline condition in our studies highlights that the absence of manipulations does not reduce uncertainty. In Studies II and III, uncertainty was lowest in manipulation conditions and highest in baseline conditions without manipulations. Across studies, detecting (or believing to detect) an alteration was consistently associated with lower credibility. The negative correlations were small to moderate in Studies I and II and largest in Study III, indicating that perceived alteration reduced credibility regardless of whether manipulations were actually present.
Additional Measures
Our findings are in line with the association between source vividness and credibility as discussed by Lee and Shin in the context of deepfakes and misinformation Lee and Shin (2022). We find, however, our source vividness measure to be more affected than our message credibility measure. This aligns with the understanding of video credibility assessment as complex and affected by multiple factors. While all participants learned from exposure to the stimulus, none of the employed manipulations affected their learning success. All learning-relevant information was mediated via the audio track of the videos. This suggests that the audio-based manipulations employed in Study II and Study III did not interfere with comprehensibility and that the measured reduction in source vividness and message credibility did not coincide with reduced apprehension.
Study Manipulation Key findings Link to others
I Visual distortion of a real recorded video. Distortion sharply reduced processing fluency and message credibility, despite unchanged video content. Establishes that non-probative visual quality alone can lower credibility, motivating the broader manipulation space in II and III.
II Reduced visual clarity, reduced speech clarity, and audio–visual asynchrony. Visual degradation and AV asynchrony lowered credibility and source vividness more than mild speech distortion; learning outcomes were unchanged. Shows that visual and timing cues systematically affect credibility, and that these effects occur without harming learning; provides the bridge to ecologically realistic deepfake cues in III.
III Synthetic face and/or synthetic narration (deepfake video). Fully synthetic face+voice reduced credibility and vividness most strongly. Visual deepfake artifacts drove larger credibility drops than synthetic narration; fluency and learning remained stable. Demonstrates that the credibility penalties from I and II generalize to realistic deepfake artifacts, with strongest effects when both modalities are synthetic. Together, the studies trace a progression from abstract distortions to real-world deepfake cues.
Table 1: Summary of key findings across Studies I–III on how visual distortions, audio distortions, and deepfake artifacts shape video credibility.
Beyond their substantive results, the studies also offer a methodological contribution for HCI. By experimentally decomposing synthetic-media manipulations into visual, auditory, and temporal components, and by separating deepfake-generated faces from deepfake-generated narration, we provide a reusable framework for probing how specific perceptual cues shape credibility judgments during interaction. This level of control is valuable for designing and evaluating transparency features, warning cues, provenance indicators, and other interface mechanisms that must target particular classes of artifacts rather than treat “deepfakes” as a single category. The open release of all stimuli, code, and preregistrations further supports iterative design and enables HCI researchers to build on well-specified, reproducible manipulations rather than ad-hoc media examples.
6.1 Limitations
This work has several limitations that qualify the scope of its claims. First, our samples were recruited via Prolific. Such panels tend to over represent relatively internet‐savvy participants. This may shift sensitivity to visual and audio irregularities compared to the general population. The direction of any bias is not obvious a priori: greater familiarity may heighten vigilance (increasing detection and decreasing credibility) or normalize minor artifacts (attenuating effects). Second, the studies were limited to U.S.-based, English-speaking participants. Cultural norms, media ecologies, and baseline trust in audiovisual media vary across societies; so may the salience of particular artifact types (e.g., eye contact norms, dubbing conventions). Our findings therefore warrant cross-cultural replications and extensions in other languages and cultures. Third, our manipulations necessarily cover only a small slice of the space of possible distortions and deepfake-generation pipelines. Different generation procedures likely produce different artifacts, and they evolve rapidly. We targeted representative classes of distortions, but did not and cannot comprehensively sample the design space. A further limitation is that each condition relied on a single video stimulus; this constrains generalizability and underscores the need for follow-up work using multiple stimuli. Beyond these core constraints, ecological factors also limit generalizability. We intentionally used an unknown presenter and an educational talk-to-camera format to minimize confounds such as source reputation and political identity; effects could differ in news, entertainment, or highly partisan contexts where social cues dominate. Finally, our key metacognitive measure, the single-slider self-reports of processing fluency, may be coarse or susceptible to response heuristics, and alteration detection was measured with a single item. More sensitive instrumentation (e.g., response latencies, secondary-task load, eye movements, or psychophysiological indices) would help triangulate the underlying processes.
6.2 Implications
Our findings suggest, media literacy curricula should explicitly train attention to video-internal cues that our studies implicate as credibility-relevant: (i) local inconsistencies in shading/illumination and skin texture; (ii) facial geometry and boundary seams; and (iii) audio–visual timing mismatches and mouth–speech incongruence. Instruction should emphasize two guardrails: absence of visible artifacts is not evidence of authenticity, and artifact detection should be combined with source and context checks rather than used in isolation.
7 Conclusion
As deepfakes proliferate, it becomes more pressing to understand how humans process manipulated videos. Understanding how video production and synthesis relate to credibility evaluations, judgments of truth, and information recall can bolster us against the malicious spread of misinformation. The present paper takes a first stab at this and investigates how video aspects other than the language-based content affect their credibility. Through three studies, we show that distortions of the visual content of videos and deepfake artifacts can negatively affect video credibility. The reduced credibility effect can be observed across manipulations that may be interpreted as indicators of deceit, such as deepfake-synthesis suitable for impersonation, and distortions that are not associated with malicious intentions, like the reduced visual clarity manipulation seen in Study I and Study II. We did not find evidence to support the hypothesis that the relationship between a video’s non-informational content and its credibility is based on processing fluency effects. More research of the cognitive processes involved in video perception and evaluation is needed, not least to assess the threat posed by deepfake videos. Our research marks a step forward in the nuanced understanding of video credibility in the age when online videos become an ever more central medium while the proliferation of deepfake technology undermines its veracity.
8 Acknowledgments
This research was supported by a Foundational Integrity Research award by Meta Research. In service of readability and conciseness, ChatGPT was utilized to copy-edit author generated content of this manuscript.
Appendix A Appendix
A.1 Appendix 1
Power Analysis
For Study II and III, a power analysis conducted with G*Power (ANOVA: fixed effects, omnibus, one-way; $f=0.22$, $\alpha=0.05$, power = $0.95$, number of groups = $4$) yielded a target sample size of $356$ participants. We recruited 400 participants to account for invalid responses.
A.2 Appendix 2
The stimulus source.
The actor was filmed in a resolution of $3840\times 2160$ in front of a green-screen that was later replaced by a gray background. The resulting video was resized to $1280\times 720$ pixels and compressed with variable bitrate (target = $0.5\text{\,}\mathrm{Mbit}\text{\,}{\mathrm{s}}^{-1}$; max = $5\text{\,}\mathrm{Mbit}\text{\,}{\mathrm{s}}^{-1}$). All editing of the stimulus source video has been carried out using the visual effects software Adobe After Effects 23.2.1.
A.3 Appendix 3
Audio-Echo Effect.
The echo effect was created with the built-in delay effect of the visual effects software Adobe After Effects 23.2.1 (delay time: $0.5\text{\,}\mathrm{s}$, delay amount: 25%, feedback: 25%, dry out: 25%, wet out: 50%).
A.4 Appendix 4
Dependent Variable Instrument Type Prompt
Subjective Processing Fluency Single item scale adapted from Graf et al. [2018] Slider anchored with “difficult” and “easy” “The process of watching this video was…”
Message Credibility Message credibility scale adapted from Appelman & Sundar Appelman and Sundar [2016] 7-point Likert-type items anchored with “describes very poorly” and “describes very well” How well does the word ‘accurate’ describe the video?
How well does the word ‘authentic’ describe the video?
How well does the word ‘believable’ describe the video?
Digital Manipulation Detection Multiple choice item adapted from Nightingale et al. [2017] Multiple choice question with the answer options: “Yes”, “I don’t know”, and “No” “Do you think the video has been digitally altered?”
Table S1: Dependent measures used in Study I.
Dependent Variable Instrument Type
Conceptual Learning Conceptual learning scale adapted from Petersen et al. [2021] Eight multiple choice questions, with four answer options each, applied before and after stimulus presentation
Factual Learning Factual learning scale adapted from Petersen et al. [2021] Eight multiple choice questions, with four answer options each, applied before and after stimulus presentation
Subjective Processing Fluency Single item scale adapted from Graf et al. [2018] Slider anchored with “difficult” and “easy”
Message Credibility Message credibility scale adapted from Appelman and Sundar [2016] Three 7-point Likert-type items anchored with “describes very poorly” and “describes very well”
Source Vividness Source Vividness scale adapted from Lee and Shin [2022] Four 7-point Likert-type items anchored with “strongly disagree” and “strongly agree”
Liking Online news liking scale adapted from Sundar [1999] Five 7-point Likert-type items anchored with “describes very poorly” and “describes very well”
Attentional Focus Attentional Focus scale adapted from Busselle and Bilandzic [2009] Three 7-point Likert-type items anchored with “strongly disagree” and “strongly agree”
Digital Manipulation Detection Multiple choice item adapted from Nightingale et al. [2017] Multiple choice question with the answer options: “Yes”, “I don’t know”, and “No”
Table S2: Dependent measures employed in Study II and Study III.
A.5 Appendix 5: Frequentist analysis
Test p-value Cohen’s d
Processing Fluency < 0.0001 1.06
Table S3: Study I: Summary of T-test Results for Processing Fluency
Test p-value Cohen’s d
Processing Fluency = 0.0006 0.38
Table S4: Study I: Summary of T-test Results for Message Credibility
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 7458 2486.2 2.844 0.0377*
Residuals 351 306806 874.1
Table S5: Study II: ANOVA Results for Processing Fluency Across Conditions
Comparison P-value
Reduced Visual Clarity vs. Baseline 0.35
Audi-Visual Incoherence vs. Baseline 0.34
Reduced Auditive Clarity vs. Baseline 1.00
Reduced Auditive Clarity vs. Reduced Visual Clarity 0.17
Audi-Visual Incoherence vs. Reduced Visual Clarity 1.00
Audi-Visual Incoherence vs. Reduced Auditive Clarity 0.17
Table S6: Study II: Pairwise Comparisons for Processing Fluency Across Conditions. P-values adjusted using Bonferroni method
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 10.3 3.450 2.954 0.0326*
Residuals 351 410.0 1.168
Table S7: Study II: ANOVA Results for Message Credibility Across Conditions
Comparison P-value
Reduced Visual Clarity vs. Baseline 0.677
Audi-Visual Incoherence vs. Baseline 0.313
Reduced Auditive Clarity vs. Baseline 1.000
Reduced Auditive Clarity vs. Reduced Visual Clarity 0.190
Audi-Visual Incoherence vs. Reduced Visual Clarity 1.000
Audi-Visual Incoherence vs. Reduced Auditive Clarity 0.075
Table S8: Study II: Pairwise Comparisons for Message Credibility Across Conditions (P-values adjusted using Bonferroni method)
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 21 6.992 1.694 0.168
Residuals 351 1448 4.126
Table S9: Study II: ANOVA Results for Factual Learning Across Conditions
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 4.6 1.55 0.787 0.502
Residuals 351 691.3 1.97
Table S10: Study II: ANOVA Results for Conceptual Learning Across Conditions
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 36.0 11.998 6.478 0.000281***
Residuals 351 650.1 1.852
Table S11: Study II: ANOVA Results for Condition on Source Vividness
Comparison P-value
Reduced Visual Clarity vs. Baseline 0.0066
Audi-Visual Incoherence vs. Baseline 0.0040
Reduced Auditive Clarity vs. Baseline 1.0000
Reduced Auditive Clarity vs. Reduced Visual Clarity 0.0372
Audi-Visual Incoherence vs. Reduced Visual Clarity 1.0000
Audi-Visual Incoherence vs. Reduced Auditive Clarity 0.0239
Table S12: Study II: Pairwise Comparisons for Source Vividness Across Conditions (P-values adjusted using Bonferroni method)
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 10.03 3.344 3.761 0.0111*
Residuals 351 312.09 0.889
Table S13: Study II: ANOVA Results for Condition on Liking
Comparison P-value
Reduced Visual Clarity vs. Baseline 0.151
Audi-Visual Incoherence vs. Baseline 0.031
Reduced Auditive Clarity vs. Baseline 1.000
Reduced Auditive Clarity vs. Reduced Visual Clarity 0.414
Audi-Visual Incoherence vs. Reduced Visual Clarity 1.000
Audi-Visual Incoherence vs. Reduced Auditive Clarity 0.102
Table S14: Study II: Pairwise Comparisons for Liking Across Conditions (P-values adjusted using Bonferroni method)
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 8.9 2.976 1.046 0.372
Residuals 351 998.3 2.844
Table S15: Study II: ANOVA Results for Condition on Attentional Focus
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 1679 559.7 0.823 0.482
Residuals 361 245374 679.7
Table S16: Study III: ANOVA Results for Processing Fluency Across Conditions
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 15.1 5.040 4.229 0.00589**
Residuals 361 430.2 1.192
Table S17: Study III: ANOVA Results for Message Credibility Across Conditions
Comparison P-value
RecVid-SynAud vs. RecVid-RecAud 0.0582
SynVid-RecAud vs. RecVid-RecAud 0.2727
SynVid-SynAud vs. RecVid-RecAud 0.0044
SynVid-RecAud vs. RecVid-SynAud 1.0000
SynVid-SynAud vs. RecVid-SynAud 1.0000
SynVid-SynAud vs. SynVid-RecAud 1.0000
Table S18: Study III: Pairwise Comparisons for Message Credibility Across Conditions (P-values adjusted using Bonferroni method)
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 40.3 13.424 6.739 0.000196***
Residuals 361 719.1 1.992
Table S19: Study III: ANOVA Results for Source Vividness Across Conditions
Comparison P-value
RecVid-SynAud vs. RecVid-RecAud 1.0000
SynVid-RecAud vs. RecVid-RecAud 0.0032
SynVid-SynAud vs. RecVid-RecAud 0.0015
SynVid-RecAud vs. RecVid-SynAud 0.0797
SynVid-SynAud vs. RecVid-SynAud 0.0458
SynVid-SynAud vs. SynVid-RecAud 1.0000
Table S20: Study III: Pairwise Comparisons for Source Vividness Across Conditions (P-values adjusted using Bonferroni method)
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 5.45 1.8161 2.183 0.0896
Residuals 361 300.26 0.8318
Table S21: Study III: ANOVA Results for Liking Across Conditions
Source Df Sum Sq Mean Sq F value Pr(>F)
Condition 3 2.5 0.8494 0.314 0.815
Residuals 361 977.3 2.7072
Table S22: Study III: ANOVA Results for Attentional Focus Across Conditions
A.6 Appendix 6: Bayesian analysis
Figure S1: Distribution of credible differences between the mean processing fluency of the baseline video and the synthesized narration video.
Figure S2: Distribution of credible differences between the mean processing fluency of the baseline video and the synthesized video.
Figure S3: Distribution of credible differences between the mean message credibility of the baseline video and the synthesized narration video.
Figure S4: Distribution of credible differences between the mean message credibility of the baseline video and the synthesized video.
Figure S5: Distribution of credible differences between the mean correct answers (Learning) of the baseline video and synthesized video and narration condition.
Figure S6: Distribution of credible differences between the mean correct answers (Learning) of the baseline video and synthesized video condition.
Figure S7: Distribution of credible differences between the mean correct answers (Learning) of the baseline video and synthesized narration condition.
Figure S8: Distribution of credible differences between the mean source vividness of the baseline video and synthesized video and narration condition.
Figure S9: Distribution of credible differences between the mean source vividness of the baseline video and synthesized video condition.
Figure S10: Distribution of credible differences between the mean source vividness of the baseline video and synthesized narration condition.
Figure S11: Distribution of credible differences between the mean liking measurement of the baseline video and synthesized video and narration condition.
Figure S12: Distribution of credible differences between the mean liking measurement of the baseline video and synthesized video condition.
Figure S13: Distribution of credible differences between the mean liking measurement of the baseline video and synthesized narration condition.
Figure S14: Distribution of credible differences between the mean attentional focus of the baseline video and synthesized video and narration condition.
Figure S15: Distribution of credible differences between the mean attentional focus of the baseline video and synthesized video condition.
Figure S16: Distribution of credible differences between the mean attentional focus of the baseline video and synthesized narration condition.
A.7 Appendix 7
Association between alteration detection and credibility
Across all three studies, we tested whether participants who believed that the video had been digitally altered rated its credibility lower. In each study, credibility differed significantly across the three alteration-detection categories (Yes, No, I don’t know) as indicated by Kruskal–Wallis tests (Study 1: $\chi^{2}(2)=20.77$, $p<.001$; Study 2: $\chi^{2}(2)=25.28$, $p<.001$; Study 3: $\chi^{2}(2)=58.96$, $p<.001$). Dunn post-hoc tests consistently showed that participants selecting Yes rated the video as less credible than those selecting No or I don’t know (all adjusted $p<.001$ across studies). Collapsing responses into “Detected” versus “Not detected” confirmed these differences via Wilcoxon rank-sum tests (Study 1: $W=9778$, $p<.001$; Study 2: $W=9750.5$, $p<.001$; Study 3: $W=7999$, $p<.001$). Spearman rank correlations likewise indicated consistent negative associations between alteration detection and credibility (Study 1: $\rho=-0.25$, $p<.001$; Study 2: $\rho=-0.27$, $p<.001$; Study 3: $\rho=-0.40$, $p<.001$). Together, these analyses show a robust pattern across all studies: believing that the video was digitally altered is reliably associated with lower perceived credibility.